OpenStudy (anonymous):

Need someone to make sure my answer is correct :) Y=x2 + 8x + 16 Find the y int and the vertex

OpenStudy (anonymous):

I found that bot are (-4,0) can someone make sure this is correct?

OpenStudy (anonymous):

Both*

OpenStudy (anonymous):

yes its right

Directrix (directrix):

@coreymacg The vertex is at (-4,0) but the y-intercept is elsewhere. The y-intercept is the value of y which has an x of 0. So crank it out. y= =x^2 + 8x + 16 y = 0^2 + 8*0 + 16 y = ?

Directrix (directrix):

Did you get the y-intercept?

OpenStudy (anonymous):

yes I did ty.

Directrix (directrix):

y = 0^2 + 8*0 + 16 y = 0 + 0 + 16 = ? --> This sum is the y-intercept. @Romero00 work is incorrect.

OpenStudy (anonymous):

y=16?